an alpha particle of momentum p is bombarded on the nucleus the distance of the closest aproch is r if momentum of alpha particle is made 6 times the distance of the closest approch becomes



Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

assume that when the alpha particle has momentum 'p', it is far enough from the nucleus so that its potential energy can be considered to be zero. kinetic energy= p2/2m, at closest approach kinetic energy=0, and potential energy=p2/2m=k/r, where k is a constant which can easily be obtained from elementary electrostatics. r=2mk/p2, thus r is inversely proportional to p2,

Therefore, if momentum of alpha particle is made 6 times, distance of closest approach becomes r/36